OpenStudy (anonymous):

Solve the equation step by step. List each step, along with the property you used during that step. -5 * 1 * 11 * 4

OpenStudy (anonymous):

-220

OpenStudy (anonymous):

How did you get that

OpenStudy (anonymous):

multiplied every digit and then put a -ve sign in the end

OpenStudy (anonymous):

Solve the equation step by step. List each step, along with the property you used during that step

OpenStudy (anonymous):

Please

OpenStudy (anonymous):

1 * 11 = 11 11 * 4 = 44 44 * (-5)= -220.... Ans

OpenStudy (anonymous):

got it..?

OpenStudy (anonymous):

Now I'm clear

OpenStudy (anonymous):

Thanks

OpenStudy (anonymous):

n.p..

OpenStudy (anonymous):

1 more thing

OpenStudy (anonymous):

yeah?

OpenStudy (anonymous):

5d = 17.1

OpenStudy (anonymous):

I'm lost please show me the work

OpenStudy (anonymous):

didnt get u.. :S

OpenStudy (anonymous):

1 * 11 = 11 11 * 4 = 44 44 * (-5)= -220.... Ans This?

OpenStudy (anonymous):

Yes I got that one

OpenStudy (anonymous):

hmm :)

OpenStudy (anonymous):

u seems a school student? right?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

I am trying to learn ahead of time for extra credit

OpenStudy (anonymous):

wow.. nice.